Evaluate the limit  "\displaystyle{\lim\limits_{x \to 0} \dfrac{e^{x}-x-1}{x^2}}" by using l'Hopital's Rule twice.




1
Expert's answer
2021-12-14T05:54:29-0500

"\\text{Evaluate the limit } \\mathop {\\lim }\\limits_{x \\to 0} \\frac{{{e^x} - x - 1}}{{{x^2}}} \\text{ by using l'Hopital's Rule twice}\n\n\\\\\n\n\\textit{Solution}\n\\\\\n\\text{by direct substitution, } \\frac{{{e^x} - x - 1}}{{{x^2}}} = \\frac{0}{0} \\text{ which is an } \\textit{Intermediate form}.\n\\\\\n\\text{Applying the l'Hopital's rule,}\n\t\\implies \\mathop {\\lim }\\limits_{x \\to 0} \\frac{{{e^x} - x - 1}}{{{x^2}}} = \\mathop {\\lim }\\limits_{x \\to 0} \\frac{{\\frac{d}{{dx}}\\left( {{e^x} - x - 1} \\right)}}{{\\frac{d}{{dx}}\\left( {{x^2}} \\right)}} = \\mathop {\\lim }\\limits_{x \\to 0} \\frac{{{e^x} - 1}}{{2x}}\n\\\\\n\\text{Again, by direct substitution, } \\mathop {\\lim }\\limits_{x \\to 0} \\frac{{{e^x} - 1}}{{2x}} = \\frac{0}{0}\n\\text{Applying the l'Hopital's rule again,}\n\t\\implies \\mathop {\\lim }\\limits_{x \\to 0} \\frac{{{e^x} - 1}}{{2x}} = \\mathop {\\lim }\\limits_{x \\to 0} \\frac{{\\frac{d}{{dx}}\\left( {{e^x} - 1} \\right)}}{{\\frac{d}{{dx}}\\left( {2x} \\right)}} = \\mathop {\\lim }\\limits_{x \\to 0} \\frac{{{e^x}}}{2} = \\frac{1}{2}\n\\\\\n\n\\text{Hence } \\mathop {\\lim }\\limits_{x \\to 0} \\frac{{{e^x} - x - 1}}{{{x^2}}} = \\frac{1}{2}"